Let f : R → R be defined by f(x) = 3√(1 – x 3 ). a. Prove that f is bijective b. Determine f -1 (x)


1
Expert's answer
2022-05-24T23:27:08-0400

a. Let "f(x_1)=f(x_2)." It means that


"\sqrt[3]{1-x_1^3}=\sqrt[3]{1-x_2^3}"

"1-x_1^3=1-x_2^3"

"(x_1-x_2)(x_1^2+x_1x_2+x_3^2)=0"




"x_1-x_2=0"

"x_1=x_2"

The function "f(x)=\sqrt[3]{1-x^3}" is bijective (one-to-one ) from "\R" to "\R."



"f(x)=\sqrt[3]{1-x^3}, x\in \R"

"y=\sqrt[3]{1-x^3}"

Change "x" and "y"


"x=\sqrt[3]{1-y^3}"

Solve for "y"


"y^3=1-x^3"




"y=\sqrt[3]{1-x^3}"

Then


"f^{-1}(x)=\sqrt[3]{1-x^3}"